The common stock of Margot, Inc. is selling for $65 a share. The par value per share is $1. Currently, the firm has a total market value of $93317. How many shares of stock will be outstanding if the firm does a 5-for-2 stock split? (Round answer to 0 decimal places, do not round intermediate calculations)
Answer #1

Number of shares currently=(93317/65)=1435.646154

Hence shares of stock after the stock-split=(1435.646154*5/2)

which is equal to

=3589 shares(Approx).